Heavy snowfall at Cerro Pachon has completely wiped out the Zorro run originally scheduled between July 11 and 17th. We expect the telescope can be operating again on Monday 18th, and the Zorro run has been extended until Friday 22nd.
After a successful realignment of its optics, the first Zorro run of the semester happened between Sept 16 and 22. Zorro has both cameras operative and waiting for proposals for 22A!
Zorro will be removed on Sept 1st for its optics realignment between Sept 8 to 10th. It will be installed back on the telescope between Sept 13 and 15th, and be ready before the Zorro run starting on the 16th.
The Zorro blue filter wheel, that had been faulty since last March, was succesfully replaced last week. An unfortunate side-effect of the procedure is that the red and blue cameras are now misaligned by about 6 arcsecods, a shift bigger than the speckle field-of-view, which makes simultaneous observations with both cameras in speckle mode impossible. Both cameras are working fine, only independently. New PIs should therefore indicate which camera they want to use. Both cameras can still be used, but this will imply doubling the requested time.
After a minor fix to its controlling computer, Zorro is back online starting May 31.
A failure in the Zorro PC will make the instrument unavailable until May ~5th.
The blue filter wheel of Zorro has been behaving erratically for past few monhts. This has had only minimal impact on operations, since the great majority of Zorro programs use the same filter. A new filter wheel has arrived recently to Pachon and the replacement will happen in the next few weeks.